OpenStudy (anonymous):

8-11). Simplify the following radical expression as much as possible. Assume all variables represent positive numbers. 2 sqrt(45 x y^2)

OpenStudy (anonymous):

\[2*\sqrt{45xy^2} = 2*\sqrt{45}*\sqrt{x}*\sqrt{y^2}= 2*\sqrt{9*5}*\sqrt{x}*y = 2y*\sqrt{9}*\sqrt{5x}=6y \sqrt{5x}\]

OpenStudy (anonymous):

it looks like it is cut off. the answer is \[6y \sqrt{5x}\]